# Addressing outliers in mixed-effects logistic regression: a more robust modeling approach

### v1

@@ Acknowledgements
The parameter $\sigma$ of the binomial-logit-$t$ and binomial-logit-normal models is assigned a half-$t$ prior with a scale of $4$. This is specified as
@@
We specify the MGH-$t$ prior distribution of Huang and Wand [14] for each regression model’s variance-covariance matrix $\bm{\Sigma}$ as a more appropriate alternative to the conventional inverse Wishart distribution. The MGH-$t$ prior distribution of $\bm{\Sigma}$ is expressed as a mixture representation of gamma distributions for the diagonal entries of the diagonal matrix $\bm{\Psi}=\text{diag}\left(\psi_{1},\ldots,\psi_{x},\ldots,\psi_{d}\right)$, and a Wishart distribution for $\bm{\Sigma}^{-1}$ with inverse scale matrix $4\bm{\Psi}$ and degrees of freedom $d+1$ [8]. The gamma distribution parameters for $\psi_{x}$ ($x=1,\ldots,d$) are set to 0.5 for the shape and $\nicefrac{{1}}{{2500}}$ for the rate, configuring the diagonal entries of $\bm{\Psi}$. This configuration results in the specification of the half-$t$ prior distribution with a location parameter of 0, a scale parameter of $2500$, and two degrees of freedom for the standard deviation terms in $\bm{\Sigma}$ and a uniform prior distribution, namely $\text{Uniform}\left(-1,1\right)$, for the correlation terms.
@@
Let $P\left(\bm{\Theta}\left|\bm{y}\right.\right)$ represent the posterior distribution of $\bm{\Theta}$ for all $\bm{y}$ (complete dataset), and $P\left(\bm{\Theta}\left|\bm{y}_{\left[ij\right]}\right.\right)$ the posterior distribution of $\bm{\Theta}$ with observation $y_{ij}$ excluded. The Monte Carlo estimate of the K-L divergence between $P\left(\bm{\Theta}\left|\bm{y}\right.\right)$ and $P\left(\bm{\Theta}\left|\bm{y}_{\left[ij\right]}\right.\right)$ under regression model $R$ is given by:
$\text{KL}_{R}\left(P\left(\bm{\Theta}\left|\bm{y}\right.\right),P\left(\bm{\Theta}\left|\bm{y}_{\left[ij\right]}\right.\right)\right)$ $=\log\left\{\frac{1}{K}\sum_{k=1}^{K}\left[P\left(y_{ij}\left|\bm{\alpha}^{\left(k\right)},\bm{u}_{i}^{\left(k\right)}\right.\right)\right]^{-1}\right\}$ (22)
$+\frac{1}{K}\sum_{k=1}^{K}\log\left[P\left(y_{ij}\left|\bm{\alpha}^{\left(k\right)},\bm{u}_{i}^{\left(k\right)}\right.\right)\right].$ (23)
We calculate the Kullback-Leibler (K-L) divergence estimates for each observation $i$ and $j$ to determine whether $y_{ij}$ under regression model $R$ is influential, effectively identifying data points that significantly affect the parameter estimates. Following the approach of Tomazella et al. [22], we consider $y_{ij}$ influential if
$0.5\left(1+\sqrt{1-\exp\left[-2\text{KL}_{R}\left(P\left(\bm{\Theta}\left|\bm{y}\right.\right),P\left(\bm{\Theta}\left|\bm{y}_{\left[ij\right]}\right.\right)\right)\right]}\right)\geq 0.99$ (24)
@@
$\beta_{\text{time}}$ $-0.07$ [-0.11; $-0.02$] $-0.02$ [-0.06; $0.01$] $-0.03$ [-0.07; $0.01$] $-0.05$ [-0.08; $-0.01$]
@@
WEB FIGURE 2: Residual diagnostics across four different models: binomial-logit-$t$, binomial-logit-normal, beta-binomial, and binomial. The QQ plots display the uniformity of the scaled residuals; a closer alignment to the 45-degree red line indicates a better fit. The binomial-logit-$t$ model shows the best fit, with residuals closely aligned across the expected range. The binomial model exhibits the most significant deviations, indicating potential issues with the model’s fit and overdispersion. The beta-binomial and binomial-logit-normal models show intermediate performance, with slight deviations.
@@ E.1 Model performance simulation study
We consider four scenarios to evaluate model performance under different conditions of overdispersion and outliers: (i) $\sigma=5$ $\nu=5$, (ii) $\sigma=5$ $\nu=30$, (iii) $\sigma=2$ $\nu=5$, and (iv) $\sigma=2$ $\nu=30$. Each of these scenarios is evaluated for two different values of the treatment effect parameter: $\beta_{\text{tx}}=0.1$ and $\beta_{\text{tx}}=0.3$. Additionally, for each combination of scenario and $\beta_{\text{tx}}$ value, we consider two different numbers of clusters: 100 clusters ($I=100$) and 300 clusters ($I=300$). For all scenarios, the remaining parameters are set as $\sigma_{a_{0}}=0.4$, $\sigma_{a_{1}}=0.3$, $\rho_{a_{0},a_{1}}=0.5$, $\beta_{0}=1.5$, and $\beta_{\text{time}}=-0.05$. Web Table 2 summarizes all 16 parameter combinations.
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0017$ $0.953$
@@
2 $\beta_{0}$ $1.50$ $-0.0164$ $0.927$
$\beta_{\text{time}}$ $-0.05$ $0.0010$ $0.951$
$\beta_{\text{tx}}$ $0.10$ $-0.0020$ $0.960$
@@
3 $\beta_{0}$ $1.50$ $-0.0175$ $0.958$
$\beta_{\text{time}}$ $-0.05$ $0.0015$ $0.946$
$\beta_{\text{tx}}$ $0.10$ $-0.0023$ $0.953$
@@
4 $\beta_{0}$ $1.50$ $-0.0344$ $0.949$
$\beta_{\text{time}}$ $-0.05$ $0.0005$ $0.941$
$\beta_{\text{tx}}$ $0.10$ $-0.0019$ $0.946$
@@
$\beta_{\text{time}}$ $-0.05$ $0.0009$ $0.945$
$\beta_{\text{tx}}$ $0.30$ $-0.0016$ $0.946$
$\beta_{\text{cnt}}$ $0.02$ $-0.0001$ $0.941$
6 $\beta_{0}$ $1.50$ $-0.0074$ $0.950$
$\beta_{\text{time}}$ $-0.05$ $-0.0004$ $0.943$
@@
7 $\beta_{0}$ $1.50$ $-0.0047$ $0.955$
$\beta_{\text{time}}$ $-0.05$ $0.0001$ $0.959$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0017$ $0.951$
@@
$\beta_{\text{cnt}}$ $0.02$ $-0.0004$ $0.948$
9 $\beta_{0}$ $1.50$ $-0.0079$ $0.951$
$\beta_{\text{time}}$ $-0.05$ $0.0004$ $0.957$
$\beta_{\text{tx}}$ $0.10$ $-0.0009$ $0.950$
@@
10 $\beta_{0}$ $1.50$ $-0.0020$ $0.941$
$\beta_{\text{time}}$ $-0.05$ $-0.0010$ $0.957$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0006$ $0.958$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0006$ $0.952$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0007$ $0.943$
@@
$\beta_{\text{time}}$ $-0.05$ $0.0012$ $0.940$
$\beta_{\text{tx}}$ $0.30$ $-0.0004$ $0.949$
$\beta_{\text{cnt}}$ $0.02$ $-0.0001$ $0.945$
15 $\beta_{0}$ $1.50$ $-0.0070$ $0.952$
$\beta_{\text{time}}$ $-0.05$ $0.0005$ $0.937$
@@
16 $\beta_{0}$ $1.50$ $-0.0112$ $0.950$
$\beta_{\text{time}}$ $-0.05$ $-0.0004$ $0.947$
$\beta_{\text{tx}}$ $0.30$ $-0.0002$ $0.954$
@@ E.2 Contamination simulation study
Contamination was introduced by first generating the response variable $y_{ij}$ using the binomial-logit-normal model and then randomly selecting a proportion of these responses based on the contamination rate; selected responses were replaced with extreme values sampled from either the lower tail (values 0, 1, 2) or the upper tail (values $m-2$, $m-1$, $m$) of the binomial-logit-normal distribution, with the same rate applied to both tails, resulting in a mixture of the binomial-logit-normal distribution and two uniform distributions.
The contamination simulation study used the following parameters: contamination rates $\in\{0,0.02,0.10\}$, $\sigma=5$, $\sigma_{a_{0}}=0.4$, $\sigma_{a_{1}}=0.3$, $\rho_{a_{0},a_{1}}=-0.5$, $\beta_{0}=1.5$, $\beta_{\text{time}}=-0.05$, $\beta_{\text{tx}}=0.1$, $\beta_{\text{cnt}}=0.02$, $I=50$, and $m=30$. We used the same data structure as in the model performance simulation study, with $J_{i}=11$ observations per cluster, where $j=1,\ldots,11$ and $\text{time}_{ij}=j-1$, and the covariates $\text{cov}_{i}$ and $\text{group}_{i}$ were simulated using the same approach.
@@
The simulation results presented in Web LABEL:tab:sim_contamination demonstrate that the binomial-logit-$t$ model outperforms the binomial-logit-normal model in the presence of data contamination, particularly in terms of RMSE and the length of the 95% HPD interval. The binomial-logit-$t$ model exhibited smaller RMSE and shorter HPD intervals, indicating more precise estimates, especially in scenarios with higher contamination rates. Bias and coverage were similar between the two models across all scenarios, consistent with the fact that the scenarios resulted in a symmetrical distribution, minimizing the potential for serious bias issues under contamination. Notably, the shorter HPD intervals with the binomial-logit-$t$ model suggest a gain in statistical power. On the other hand, the binomial-logit-normal model showed greater sensitivity to outliers, with larger RMSE and wider HPD intervals under higher contamination levels. These results suggest that while both models perform comparably in terms of bias and coverage, the binomial-logit-$t$ model offers greater robustness, precision, and statistical power in contaminated data scenarios.
@@
0% Binomial-logit-normal $\beta_{0}$ $1.50$ $-0.0102$ $0.7378$ $3.6752$ $0.947$
$\beta_{\text{time}}$ $-0.05$ $0.0013$ $0.0489$ $0.2414$ $0.950$
@@
Binomial-logit-$t$ $\beta_{0}$ $1.50$ $-0.0129$ $0.7475$ $3.6635$ $0.941$
$\beta_{\text{time}}$ $-0.05$ $0.0011$ $0.0486$ $0.2410$ $0.952$
$\beta_{\text{tx}}$ $0.10$ $-0.0001$ $0.0622$ $0.3229$ $0.959$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0088$ $0.0484$ $0.2515$ $0.949$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0019$ $0.0465$ $0.2408$ $0.951$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0164$ $0.0556$ $0.2607$ $0.936$
$\beta_{\text{tx}}$ $0.10$ $-0.0044$ $0.0681$ $0.3428$ $0.946$
@@
$\beta_{\text{time}}$ $-0.05$ $-0.0072$ $0.0487$ $0.2353$ $0.938$
$\beta_{\text{tx}}$ $0.10$ $-0.0034$ $0.0643$ $0.3161$ $0.948$

### v2

@@ Acknowledgements
Appendix. Pseudo-median vs. exact median
We show that the closed-form pseudo-median
$\widehat{y}=g\left(\eta\right)=m_{i}\text{logit}^{-1}\left(\eta\right)$ (20)
differs from the discrete median of the binomial-logit-$t$ mixture by strictly less than one count. Define
$k_{i-}=\min\left\{k_{i}\in\mathbb{Z}:P\left(Y_{i}\leq k_{i}\right)\geq 0.5\right\},$ (21)
the left-most integer whose cumulative probability reaches 50%. By definition, $P\left(Y_{i}\leq k_{i-}-1\right)<0.5$.
Since $\eta$ is the median of $\xi_{i}$ and the logistic map is strictly increasing, $\text{logit}^{-1}\left(\eta\right)$ is the median on the probability scale. However, once the binomial draw is introduced, $Y_{i}$ can only take integer values, so its median may not equal $\widehat{y}$ exactly. We analyze two possibilities, using the fact that $Y_{i}$ is stochastically increasing in $\xi_{i}$: a higher $\xi_{i}$ induces a larger success probability $p_{i}$, thus typically increasing $Y_{i}$.
Case 1:
$\widehat{y}=g\left(\eta\right)<k_{i-}.$ (22)
Then with probability above $0.5$, we must have
$\xi_{i}<g^{-1}\left(k_{i-}\right),$ (23)
pushing $Y_{i}<k_{i-}$ more than half the time and contradicting the definition of $k_{i-}$ as the first integer with $P\left(Y_{i}\leq k_{i-}\right)\geq 0.5$.
Case 2:
$\widehat{y}=g\left(\eta\right)\geq k_{i-}+1.$ (24)
Then with probability exceeding $0.5$,
$\xi_{i}\geq g^{-1}\left(k_{i-}+1\right),$ (25)
forcing $Y_{i}$ to exceed $k_{i-}$ more than half the time. Again, this contradicts the minimality of $k_{i-}$.
Since neither case can occur, we conclude
$k_{i-}\leq\widehat{y}<k_{i-}+1.$ (26)
Hence, the pseudo-median always lies between the two adjacent integers that bracket the true median, differing by strictly less than one count. Consequently, $\widehat{y}$ is a reliable surrogate for the median and is significantly easier to evaluate or interpret.
@@
The parameter $\sigma$ in both the binomial-logit-$t$ and binomial-logit-normal models is assigned a half-$t$ prior with location 0, scale $4$, and $2$ degrees of freedom. This is specified as
@@
We specify the MGH-$t$ prior distribution of Huang and Wand [17] for each regression model’s variance-covariance matrix $\bm{\Sigma}$ as a more appropriate alternative to the conventional inverse Wishart distribution. The MGH-$t$ prior distribution of $\bm{\Sigma}$ is expressed as a mixture representation of gamma distributions for the diagonal entries of the diagonal matrix $\bm{\Psi}=\text{diag}\left(\psi_{1},\ldots,\psi_{x},\ldots,\psi_{d}\right)$, and a Wishart distribution for $\bm{\Sigma}^{-1}$ with inverse scale matrix $4\bm{\Psi}$ and degrees of freedom $d+1$ [8]. The gamma distribution parameters for $\psi_{x}$ ($x=1,\ldots,d$) are set to 0.5 for the shape and $\nicefrac{{1}}{{2500}}$ for the rate, configuring the diagonal entries of $\bm{\Psi}$. This configuration results in the specification of the half-$t$ prior distribution with a location parameter of 0, a scale parameter of $2500$, and $2$ degrees of freedom for the standard deviation terms in $\bm{\Sigma}$ and a uniform prior distribution, namely $\text{Uniform}\left(-1,1\right)$, for the correlation terms.
@@
[…]
We calculate the K-L divergence estimates for each observation $i$ and $j$ to determine whether $y_{ij}$ under regression model $R$ is influential, effectively identifying data points that significantly affect the parameter estimates. Following the approach of Tomazella et al. [26], we consider $y_{ij}$ influential if
$0.5\left(1+\sqrt{1-\exp\left[-2\text{KL}_{R}\left(P\left(\bm{\Theta}\left|\bm{y}\right.\right),P\left(\bm{\Theta}\left|\bm{y}_{\left[ij\right]}\right.\right)\right)\right]}\right)\geq 0.99.$ (32)
@@
$\beta_{\text{time}}$ 0.07-0.07 [-0.11; 0.02-0.02] 0.02-0.02 [-0.06; $0.01$] 0.03-0.03 [-0.07; $0.01$] 0.05-0.05 [-0.08; 0.01-0.01]
@@
WEB FIGURE 2: Residual diagnostics across four different models: binomial-logit-$t$, binomial-logit-normal, beta-binomial, and binomial. The $Q$ plots display the uniformity of the scaled residuals; a closer alignment to the 45-degree red line indicates a better fit. The binomial-logit-$t$ model shows the best fit, with residuals closely aligned across the expected range. The binomial model exhibits the most significant deviations, indicating potential issues with the model’s fit and overdispersion. The beta-binomial and binomial-logit-normal models show intermediate performance, with slight deviations.
@@ E.1 Model performance simulation study
We consider four scenarios to evaluate model performance under different conditions of overdispersion and outliers: (i) $\sigma=5$ & $\nu=5$, (ii) $\sigma=5$ & $\nu=30$, (iii) $\sigma=2$ & $\nu=5$, and (iv) $\sigma=2$ & $\nu=30$. Each of these scenarios is evaluated for two different values of the treatment effect parameter: $\beta_{\text{tx}}=0.1$ and $\beta_{\text{tx}}=0.3$. Additionally, for each combination of scenario and $\beta_{\text{tx}}$ value, we consider two different numbers of clusters: 100 clusters ($I=100$) and 300 clusters ($I=300$). For all scenarios, the remaining parameters are set as $\sigma_{a_{0}}=0.4$, $\sigma_{a_{1}}=0.3$, $\rho_{a_{0},a_{1}}=0.5$, $\beta_{0}=1.5$, and $\beta_{\text{time}}=-0.05$. Web Table 2 summarizes all 16 parameter combinations.
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0017-0.0017 $0.953$
@@
2 $\beta_{0}$ $1.50$ 0.0164-0.0164 $0.927$
$\beta_{\text{time}}$ 0.05-0.05 $0.0010$ $0.951$
$\beta_{\text{tx}}$ $0.10$ 0.0020-0.0020 $0.960$
@@
3 $\beta_{0}$ $1.50$ 0.0175-0.0175 $0.958$
$\beta_{\text{time}}$ 0.05-0.05 $0.0015$ $0.946$
$\beta_{\text{tx}}$ $0.10$ 0.0023-0.0023 $0.953$
@@
4 $\beta_{0}$ $1.50$ 0.0344-0.0344 $0.949$
$\beta_{\text{time}}$ 0.05-0.05 $0.0005$ $0.941$
$\beta_{\text{tx}}$ $0.10$ 0.0019-0.0019 $0.946$
@@
$\beta_{\text{time}}$ 0.05-0.05 $0.0009$ $0.945$
$\beta_{\text{tx}}$ $0.30$ 0.0016-0.0016 $0.946$
$\beta_{\text{cnt}}$ $0.02$ 0.0001-0.0001 $0.941$
6 $\beta_{0}$ $1.50$ 0.0074-0.0074 $0.950$
$\beta_{\text{time}}$ 0.05-0.05 0.0004-0.0004 $0.943$
@@
7 $\beta_{0}$ $1.50$ 0.0047-0.0047 $0.955$
$\beta_{\text{time}}$ 0.05-0.05 $0.0001$ $0.959$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0017-0.0017 $0.951$
@@
$\beta_{\text{cnt}}$ $0.02$ 0.0004-0.0004 $0.948$
9 $\beta_{0}$ $1.50$ 0.0079-0.0079 $0.951$
$\beta_{\text{time}}$ 0.05-0.05 $0.0004$ $0.957$
$\beta_{\text{tx}}$ $0.10$ 0.0009-0.0009 $0.950$
@@
10 $\beta_{0}$ $1.50$ 0.0020-0.0020 $0.941$
$\beta_{\text{time}}$ 0.05-0.05 0.0010-0.0010 $0.957$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0006-0.0006 $0.958$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0006-0.0006 $0.952$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0007-0.0007 $0.943$
@@
$\beta_{\text{time}}$ 0.05-0.05 $0.0012$ $0.940$
$\beta_{\text{tx}}$ $0.30$ 0.0004-0.0004 $0.949$
$\beta_{\text{cnt}}$ $0.02$ 0.0001-0.0001 $0.945$
15 $\beta_{0}$ $1.50$ 0.0070-0.0070 $0.952$
$\beta_{\text{time}}$ 0.05-0.05 $0.0005$ $0.937$
@@
16 $\beta_{0}$ $1.50$ 0.0112-0.0112 $0.950$
$\beta_{\text{time}}$ 0.05-0.05 0.0004-0.0004 $0.947$
$\beta_{\text{tx}}$ $0.30$ 0.0002-0.0002 $0.954$
@@ E.2 Contamination simulation study
Contamination was introduced by first generating the response variable $y_{ij}$ using the binomial-logit-normal model and then randomly selecting a proportion of these responses based on the contamination rate; selected responses were replaced with extreme values sampled from either the lower tail (values 0, 1, 2) or the upper tail (values m2m-2, m1m-1, $m$) of the binomial-logit-normal distribution, with the same rate applied to both tails, resulting in a mixture of the binomial-logit-normal distribution and two uniform distributions.
The contamination simulation study used the following parameters: contamination rates $\in\left\{0,0.02,0.10\right\}$, $\sigma=5$, $\sigma_{a_{0}}=0.4$, $\sigma_{a_{1}}=0.3$, $\rho_{a_{0},a_{1}}=-0.5$, $\beta_{0}=1.5$, $\beta_{\text{time}}=-0.05$, $\beta_{\text{tx}}=0.1$, $\beta_{\text{cnt}}=0.02$, $I=50$, and $m=30$. We used the same data structure as in the model performance simulation study, with $J_{i}=11$ observations per cluster, where $j=1,\ldots,11$ and $\text{time}_{ij}=j-1$, and the covariates $\text{cov}_{i}$ and $\text{group}_{i}$ were simulated using the same approach.
@@
The simulation results presented in Web LABEL:tab:sim_contamination demonstrate that the binomial-logit-$t$ model outperforms the binomial-logit-normal model in the presence of data contamination, particularly in terms of RMSE and the length of the 95% HPD interval. The binomial-logit-$t$ model exhibited smaller RMSE and shorter HPD intervals, indicating more precise estimates, especially in scenarios with higher contamination rates. Bias and coverage were similar between the two models across all scenarios, consistent with the fact that the scenarios resulted in a symmetrical distribution, minimizing the potential for serious bias issues under contamination. Notably, the shorter HPD intervals with the binomial-logit-$t$ model suggest a gain in statistical power. On the other hand, the binomial-logit-normal model showed greater sensitivity to outliers, with larger RMSE and wider HPD intervals under higher contamination levels. These results suggest that while both models perform comparably in terms of bias and coverage, the binomial-logit-$t$ model provides greater robustness, precision, and statistical power in contaminated data scenarios.
@@
0% Binomial-logit-normal $\beta_{0}$ $1.50$ 0.0102-0.0102 $0.7378$ $3.6752$ $0.947$
$\beta_{\text{time}}$ 0.05-0.05 $0.0013$ $0.0489$ $0.2414$ $0.950$
@@
Binomial-logit-$t$ $\beta_{0}$ $1.50$ 0.0129-0.0129 $0.7475$ $3.6635$ $0.941$
$\beta_{\text{time}}$ 0.05-0.05 $0.0011$ $0.0486$ $0.2410$ $0.952$
$\beta_{\text{tx}}$ $0.10$ 0.0001-0.0001 $0.0622$ $0.3229$ $0.959$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0088-0.0088 $0.0484$ $0.2515$ $0.949$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0019-0.0019 $0.0465$ $0.2408$ $0.951$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0164-0.0164 $0.0556$ $0.2607$ $0.936$
$\beta_{\text{tx}}$ $0.10$ 0.0044-0.0044 $0.0681$ $0.3428$ $0.946$
@@
$\beta_{\text{time}}$ 0.05-0.05 0.0072-0.0072 $0.0487$ $0.2353$ $0.938$
$\beta_{\text{tx}}$ $0.10$ 0.0034-0.0034 $0.0643$ $0.3161$ $0.948$